solve the given initail value problem using laplace transformation

а - g4 22 (0) – отто) -
Answer #1

We have y = 3+2+ · 910= o=y(o) Taking Laplace on both sides, we get Lly) = b3+2+) How, Lly) = 3²4151 - sf (o)=9 (0) → s?